OpenStudy (anonymous):

prove: 1/ 1-sinx = 1+sinx/cos^2 x

OpenStudy (aravindg):

Start from RHS Replace cos^2 x by 1-sin^2 x

OpenStudy (anonymous):

it'll turn to 1+sinx/ 1-sin^2x but I don't know how it'll get to 1/1-sinx because you cant just cross out the sinx

OpenStudy (aravindg):

Yeah there you have to apply simply z^2-b^2=(a+b)(a-b) Got the idea? ;)

OpenStudy (anonymous):

huh?

OpenStudy (aravindg):

I mean a^2-b^2=(a+b)(a-b) 1-sin^2x=.........

OpenStudy (anonymous):

I dont know what youre trying to say...

OpenStudy (aravindg):

1+sin x/cos^2 x=(1+sinx)/(1-sin^2 x)=(1+sin x)/(1-sin x)(1+sin x) Can you do it now?

OpenStudy (anonymous):

ohhhh...i get it now...thanks...

OpenStudy (aravindg):

yw :)